### v1

# Unifying Quantum and Classical Dynamics

Abstract
Classical and quantum physics represent two distinct theories; however, quantum physics is regarded as the more fundamental of the two. It is posited that classical mechanics should arise from quantum mechanics under certain limiting conditions. Nevertheless, this remains a challenging objective. In this work, we explore the potential for unifying the dynamics of classical and quantum physics. This discussion does not suggest that classical behavior emerges from quantum mechanics; rather, it demonstrates the exact equivalence between the dynamics of quantum observables and their classical counterparts. It is shown that the Heisenberg equations of motion can be cast in a form that is identical to Newton’s equations of motion, with $\hbar$ being absent from the formulation. This implies that both quantum and classical dynamics are governed by the same equations, with the Heisenberg operators substituting the classical observables.
I Introduction
Quantum mechanics represents an exceptionally successful framework for elucidating the behavior of microscopic particles. This theory has not only been effectively utilized in a wide array of physical phenomena, but it has also undergone rigorous testing through precise experimental methods. Despite being primarily developed through intuition rather than strict derivation, it has demonstrated remarkable accuracy and has not required any modifications or adjustments. Given its effective description of micro objects, there is a prevailing belief that it should similarly be capable of describing macro objects, which are after all composed of these microscopic particles. The dynamics of macro objects, as is common knowledge, is described well by classical mechanics, or more familiarly, by Newton’s laws of motion. This naturally leads one to the expectation that classical mechanics should emerge out of quantum mechanics. The mechanism by which it may happen is under active debate with widely divergent views. At the most modest end is the correspondence principle proposed by Niels Bohr CP1 , which may be interpreted as classical behavior of quantum systems emerging in the limit of large quantum numbers. Another approach is the Wentzel–Kramers–Brillouin approximation, or the WKB method in which one imagines a semi-classical scenario where $\hbar$ is considered a small parameter WKB . In this approach, classical regime would correspond to assuming $\hbar$ to be negligibly small. Similarly, in the Feynman path-integral formulation of quantum mechanics FeynmanHibbs , the classical principle of least action emerges in the $\hbar\to 0$ limit due to the rapid oscillation of the phase $e^{iS/\hbar}$, which suppresses non-stationary paths and selects trajectories satisfying $\delta S=0$, corresponding to the classical path. At the radical end are ideas that quantum theory needs nonlinear modification in order to fully recover classical mechanics collapse ; gravity . Somewhere in between are ideas like decoherence where classical behavior is an outcome of weak interaction of quantum systems with certain environment degrees of freedom dec .
Before we are anywhere near answering the above questions, we may try to explore if there is some common ground between the two theories. Identifying such a common ground may be helpful in finally finding the solution to the problem. More specifically we address the issue of dynamics of observable quantities in the two theories. In classical mechanics the dynamics of observable quantities is governed by Newton’s equation of motion. In quantum mechanics there are two distinct ways to study the dynamics. One is via the Schrödinger equation which governs the time evolution of the quantum state, more commonly called the wavefunction. The other is via the Heisenberg equation of motion which governs the time evolution of operators which represent the observable quantities. Since the wavefunction doesn’t have a classical counterpart, it seems natural to look for a connection between the Heisenberg equation of motion and the Newton’s equation.
II Heisenberg picture
To start with the simplest case, we consider a particle of mass $m$, moving in one dimension, in a potential $V$. The Hamiltonian, which is the operator for energy, is then given by
$\boldsymbol{H}=\frac{\boldsymbol{p}^{2}}{2m}+V(\boldsymbol{x}),$ (1)
$\boldsymbol{x},\boldsymbol{p}$ being the operators for the position of momentum of the particle, respectively. The Heisenberg equation of motion for position can be written as
$\frac{d}{dt}\boldsymbol{x}(t)=\frac{\iota}{\hbar}[\boldsymbol{H},\boldsymbol{x}(t)],$ (2)
where $\boldsymbol{x}(t)=\boldsymbol{U}_{t}\boldsymbol{x}(0)\boldsymbol{U}^{\dagger}_{t}$, $\boldsymbol{U}_{t}=\exp(\iota\boldsymbol{H}t/\hbar)$ being the unitary time evolution operator. Since $\boldsymbol{U}_{t}$ commutes with $\boldsymbol{H}$, and $\boldsymbol{x}$ commutes with $V(\boldsymbol{x})$, the above equation yields
$m\frac{d}{dt}\boldsymbol{x}(t)=\boldsymbol{p}(t).$ (3)
Similarly, the Heisenberg equation of motion for the momentum operator can be written as
$\frac{d}{dt}\boldsymbol{p}(t)=\frac{\iota}{\hbar}[\boldsymbol{H},\boldsymbol{p}(t)]=\frac{\iota}{\hbar}\boldsymbol{U}_{t}[V(\boldsymbol{x}),\boldsymbol{p}(0)]\boldsymbol{U}^{\dagger}_{t}.$ (4)
Now let us assume that the potential $V$ is an analytical function of $x$ and it can be expanded in a Taylor series, say, around $x_{0}$
$V(x)=\sum_{n=0}^{\infty}\frac{V^{(n)}(x_{0})}{n!}(x-x_{0})^{n},$ (5)
where $V^{(n)}(x_{0})$ denotes the $n$th derivative of $V(x)$ at $x=x_{0}$. Using the series form of the potential, the commutator of $\boldsymbol{p}$ with the potential can be written as Sakurai
$[V(\boldsymbol{x}),\boldsymbol{p}(0)]$ $=$ $\sum_{n=0}^{\infty}\frac{V^{(n)}(x_{0})}{n!}[(\boldsymbol{x}-x_{0})^{n},\boldsymbol{p}]$ (6)
$=$ $\iota\hbar\sum_{n=0}^{\infty}\frac{V^{(n)}(x_{0})}{n!}n(\boldsymbol{x}-x_{0})^{n-1}$
$=$ $i\hbar V^{\prime}(\boldsymbol{x}),$
where we have used the well known commutation relation $\boldsymbol{x}^{n},\boldsymbol{p}]=\iota\hbar n\boldsymbol{x}^{n-1}$. Then the Heisenberg equation of motion for $\boldsymbol{p}$ (4) assumes the form
$\frac{d}{dt}\boldsymbol{p}(t)=-V^{\prime}(\boldsymbol{x}(t)).$ (7)
Using (3) and (7), one can write
$m\frac{d^{2}}{dt^{2}}\boldsymbol{x}(t)=\frac{d}{dt}\boldsymbol{p}(t)=-V^{\prime}(\boldsymbol{x}(t)).$ (8)
This is a fully quantum mechanical equation involving operators, but has the exact form of the Newton’s equation of motion. The only assumption that we made in arriving at this equation is that the potential is an analytical function which can be expanded in a Taylor series. Remarkable is the fact that although (8) is a fully quantum mechanical equation, it does not contain $\hbar$. Unlike the WKB method, one doesn’t need to make any approximation like $\hbar$ small” in order to establish the quantum-classical connection.
It may be instructive to look at a couple of specific examples. For a harmonic oscillator with potential energy $V(x)=\frac{1}{2}m\omega^{2}x^{2}$, the Heisenberg equation looks like
$m\frac{d^{2}}{dt^{2}}\boldsymbol{x}(t)=-m\omega^{2}\boldsymbol{x}(t),$ (9)
which is the same as Newton’s equation for a Harmonic oscillator. It is easy to see that generalization to three dimensions and many particles is straightforward, simply because of observables of different particles commute with each others. As an example, we focus on a small mass $m$ that is orbiting around a larger mass $M$. We assume that $M$ is significantly greater than $m$, which allows us to position the center of mass of the system at the center of the larger mass, designated as the origin. The Hamiltonian is given by
$\boldsymbol{H}=\frac{\vec{\boldsymbol{P}}^{2}}{2M}+\frac{\vec{\boldsymbol{p}}^{2}}{2m}-\frac{GmM}{|\vec{\boldsymbol{r}}|},$
where $\vec{\boldsymbol{r}}$ is the three-dimensional position operator of the lighter particle in the center-of-mass frame. Following the procedure described above, the Heisenberg equation for the momentum operator of the smaller particle is given by
$\frac{d}{dt}\vec{\boldsymbol{p}}(t)=-\frac{GmM}{\vec{\boldsymbol{r}}^{2}}\hat{r}$, (10)
where $\hat{r}$ is a unit vector in the direction of $\vec{\boldsymbol{r}}$. Again we observe that the Heisenberg equation coincides with the Newton’s equation.
III Charged Particle in a Electro-Magnetic Field
In the previous section, we analyzed the dynamics of systems whose equations of motion follow directly from a potential energy due to the velocity-independent nature of the forces involved. We now turn to systems in which the force depends explicitly on the particle velocity and examine whether the Heisenberg picture reproduces the corresponding classical equations of motion for such systems.
As an illustrative and physically motivated example, we consider the motion of a charged particle in an external electromagnetic field, where the force depends explicitly on the velocity through its coupling to the magnetic field. The Hamiltonian of the system is given by:
$\boldsymbol{H}=\frac{1}{2m}\left(\vec{\boldsymbol{p}}-q\,\vec{A}(\boldsymbol{\mathbf{r}})\right)^{2}-q\vec{\mathcal{E}}\cdot\vec{\boldsymbol{r}}.$ (11)
where $m$ and $q$ denote the mass and charge of the particle, respectively. The magnetic and electric fields are given by $\vec{B}=\vec{\nabla}\times\vec{A}(\mathbf{r})$ and $\vec{\mathcal{E}}$, with $\vec{A}(\mathbf{r})$ denoting the vector potential. The velocity operator is defined via the Heisenberg equation of motion:
$\boldsymbol{v}_{i}=\frac{d\boldsymbol{r}_{i}}{dt}=\frac{\iota}{\hbar}\left[\boldsymbol{H},\boldsymbol{r}_{i}\right],$ (12)
where the indices $i=1,2,3$ represent the $x,y,z$ component of the relevant quantity, respectively. Using the canonical commutation relations
$\boldsymbol{r}_{i},\boldsymbol{p}_{j}]=\iota\hbar\,\delta_{ij},\qquad[\boldsymbol{r}_{i},A_{j}(\boldsymbol{r})]=0,$ (13)
we compute
$[\boldsymbol{H},\boldsymbol{r}_{i}]$ $=\frac{1}{2m}\left[(\boldsymbol{p}_{j}-qA_{j})^{2},\boldsymbol{r}_{i}\right]$
$=\frac{1}{2m}\left((\boldsymbol{p}_{j}-qA_{j})[\boldsymbol{p}_{j},\boldsymbol{r}_{i}]+[\boldsymbol{p}_{j},\boldsymbol{r}_{i}](\boldsymbol{p}_{j}-qA_{j})\right)$
$=-\frac{\iota\hbar}{m}(\boldsymbol{p}_{i}-qA_{i}).$ (14)
Therefore, the velocity operator is
$\boldsymbol{v}_{i}=\frac{1}{m}\left(\boldsymbol{p}_{i}-qA_{i}\right)$ (15)
It is convenient to define the kinetic momentum operator
$\boldsymbol{\pi}_{i}\equiv\boldsymbol{p}_{i}-qA_{i},\qquad\boldsymbol{v}_{i}=\frac{\boldsymbol{\pi}_{i}}{m}.$ (16)
The Heisenberg equation for $\boldsymbol{\pi}_{i}$ reads
$\frac{d\boldsymbol{\pi}_{i}}{dt}=\frac{\iota}{\hbar}\left[\boldsymbol{H},\boldsymbol{\pi}_{i}\right].$ (17)
Since
$\boldsymbol{H}=\sum_{i=1}^{3}\left(\frac{1}{2m}\boldsymbol{\pi}_{i}\boldsymbol{\pi}_{i}-q\mathcal{E}_{i}\boldsymbol{r_{i}}\right),$ (18)
we obtain
$\boldsymbol{H},\boldsymbol{\pi}_{i}]=\frac{1}{2m}\left(\boldsymbol{\pi}_{j}[\boldsymbol{\pi}_{j},\boldsymbol{\pi}_{i}]+[\boldsymbol{\pi}_{j},\boldsymbol{\pi}_{i}]\boldsymbol{\pi}_{j}\right)-\iota\hbar q\mathcal{E}_{i}.$ (19)
Using
$\boldsymbol{p}_{i},f(\boldsymbol{r})]=-\iota\hbar\,\partial_{i}f(\boldsymbol{r}),$ (20)
we compute
$[\boldsymbol{\pi}_{i},\boldsymbol{\pi}_{j}]$ $=[\boldsymbol{p}_{i}-qA_{i},\boldsymbol{p}_{j}-qA_{j}]$
$=-q[\boldsymbol{p}_{i},A_{j}]-q[A_{i},\boldsymbol{p}_{j}]$
$=\iota\hbar q(\partial_{i}A_{j}-\partial_{j}A_{i})$
$=\iota\hbar q\sum_{k=1}^{3}\epsilon_{ijk}B_{k},$ (21)
where $\epsilon_{ijk}$ is the Levi-Civita epsilon. Substituting into the Hamiltonian commutator,
$\boldsymbol{H},\boldsymbol{\pi}_{i}]=\frac{\iota\hbar q}{m}\sum_{j,k}\epsilon_{ijk}\boldsymbol{\pi}_{j}B_{k}-\iota\hbar q\mathcal{E}_{i}.$ (22)
Therefore,
$m\frac{d\boldsymbol{v}_{i}}{dt}$ $=\frac{\iota}{\hbar}[\boldsymbol{H},\boldsymbol{\pi}_{i}]$ (23)
$=q\sum_{j,k}\epsilon_{ijk}\boldsymbol{v}_{j}B_{k}+q\mathcal{E}_{i},$ (24)
where $\sum_{j,k}\epsilon_{ijk}\boldsymbol{v}_{j}B_{k}$ is essentially $(\vec{\boldsymbol{v}}\times\vec{B})_{i}$. In vector form,
$m\frac{d\vec{\boldsymbol{v}}}{dt}=q\,\vec{\boldsymbol{v}}\times\vec{B}+q\vec{\mathcal{E}}$ (25)
Remarkably, the Lorentz force law emerges exactly from operator algebra in quantum mechanics.
IV Ehrenfest theorem
Till now we have studied the dynamics of quantum observables at the operator level. The values of observables depend on the quantum state. In an arbitrary quantum state we can always define the expectation values of all observables. If we are interested in looking at the experimentally accessible quantities, the expectation values serve the purpose. One can take the expectation value of both sides of equation (8) and get
$m\frac{d^{2}}{dt^{2}}\langle\boldsymbol{x}\rangle=\frac{d}{dt}\langle\boldsymbol{p}\rangle=-\langle V^{\prime}({\boldsymbol{x}})\rangle,$ (26)
where the angular brackets denote the expectation value. This is the familiar equation denoting Ehrefest theorem Ehrenfest . It shows that the expection values of position and momentum approximately follow Newton’s equations of motion. It is approximate because $\langle V^{\prime}({\boldsymbol{x}})\rangle\neq V^{\prime}(\langle{\boldsymbol{x}}\rangle)$.
For a general potential, if we can expand its derivative in a series, $V^{\prime}({\boldsymbol{x}})\approx\alpha\boldsymbol{x}+\beta\boldsymbol{x}^{2}$, then
$\langle V^{\prime}({\boldsymbol{x}})\rangle$ $=$ $\alpha\langle\boldsymbol{x}\rangle+\beta\langle{\boldsymbol{x}^{2}}\rangle=\alpha\langle\boldsymbol{x}\rangle+\beta\langle{\boldsymbol{x}}\rangle^{2}+\beta(\Delta x)^{2}$ (27)
$=$ $V^{\prime}(\langle{\boldsymbol{x}}\rangle)+\beta(\Delta x)^{2},$
where $\Delta x$ is the uncertainty in position in the given quantum state. So, if the state is such that the uncertainty in position is small, then $\langle V^{\prime}({\boldsymbol{x}})\rangle\approx V^{\prime}(\langle{\boldsymbol{x}}\rangle)$, and one can say that the expectation values follow classical dynamics. For a harmonic oscillator, since $V^{\prime}(\boldsymbol{x})$ is linear in $\boldsymbol{x}$, $\langle V^{\prime}({\boldsymbol{x}})\rangle=V^{\prime}(\langle{\boldsymbol{x}}\rangle)$, and the position and momentum expectation values exactly follow classical dynamics, for all quantum states.
V Conclusions
In this work, we have demonstrated that the dynamical equations governing classical and quantum mechanics possess a deeper and more exact correspondence than is usually emphasized. By focusing on the Heisenberg picture, where observable quantities evolve in time, we showed that the equations of motion for quantum operators can be written in a form identical to Newton’s equations of classical mechanics. Remarkably, this equivalence holds exactly at the operator level and does not require any limiting procedure such as large quantum numbers, small $\hbar$, or semiclassical approximations for both velocity-independent and velocity-dependent forces.
For a wide class of systems, ranging from particles moving in arbitrary analytic potentials to charged particles in external electromagnetic fields, the Heisenberg equations reproduce the familiar classical laws of motion, including Newton’s second law and the Lorentz force equation. The absence of $\hbar$ in these operator equations highlights that the distinction between classical and quantum dynamics does not arise from the form of the equations themselves, but rather from the non-commuting nature of quantum observables and the role of the quantum state.
We further clarified how this operator-level equivalence relates to experimentally accessible quantities through the Ehrenfest theorem. While expectation values of observables obey equations reminiscent of classical dynamics, deviations from classical behavior emerge due to quantum uncertainties and higher-order moments. Only under specific conditions, such as narrow wave packets or linear forces, do expectation values exactly follow classical trajectories.
Our results suggest that classical and quantum mechanics are dynamically unified at a fundamental level, with the key difference residing not in the laws of motion, but in the algebraic structure of observables and the interpretation of physical quantities. This perspective provides a transparent and exact common ground between the two theories and may offer useful insight into longstanding questions concerning the quantum-classical boundary.
Acknowledgements.
ARS acknowledges financial support from the Ministry of Minority Affairs, Government of India, via Maulana Azad National Fellowship under Grant No. F. 82-27/2019 (SA-III). TQ acknowledges useful discussions with Bibhash Paul.
Author declarations
The authors have no conflicts to disclose.

### v2

@@
 It is shown that the Heisenberg equations of motion can be cast in a form that is identical to Newton’s equations of motion, with $\hbar$ being absent from the formulation.
+In a generalized analysis, the Heisenberg equations are cast in a form that is identical to the classical Hamilton’s equations of motion.
 This implies that both quantum and classical dynamics are governed by the same equations, with the Heisenberg operators substituting the classical observables.
 I Introduction
-Quantum mechanics represents an exceptionally successful framework for elucidating the behavior of microscopic particles.
+Quantum mechanics represents an exceptionally successful framework for elucidating the behavior of micro particles.
 This theory has not only been effectively utilized in a wide array of physical phenomena, but it has also undergone rigorous testing through precise experimental methods.
 Despite being primarily developed through intuition rather than strict derivation, it has demonstrated remarkable accuracy and has not required any modifications or adjustments.
-Given its effective description of micro objects, there is a prevailing belief that it should similarly be capable of describing macro objects, which are after all composed of these microscopic particles.
+Given its effective description of micro objects, there is a prevailing belief that it should similarly be capable of describing macro objects, which are after all composed of these micro particles.
 The dynamics of macro objects, as is common knowledge, is described well by classical mechanics, or more familiarly, by Newton’s laws of motion.
@@
 Somewhere in between are ideas like decoherence where classical behavior is an outcome of weak interaction of quantum systems with certain environment degrees of freedom dec .
+There is also a diametrically opposite view according to which quantum mechanics is not a fundamental theory, but rather an emergent phenomenon adler .
 Before we are anywhere near answering the above questions, we may try to explore if there is some common ground between the two theories.
@@
 $\boldsymbol{x},\boldsymbol{p}$ being the operators for the position of momentum of the particle, respectively. The Heisenberg equation of motion for position can be written as
-$\frac{d}{dt}\boldsymbol{x}(t)=\frac{\iota}{\hbar}[\boldsymbol{H},\boldsymbol{x}(t)],$ (2)
-where $\boldsymbol{x}(t)=\boldsymbol{U}_{t}\boldsymbol{x}(0)\boldsymbol{U}^{\dagger}_{t}$, $\boldsymbol{U}_{t}=\exp(\iota\boldsymbol{H}t/\hbar)$ being the unitary time evolution operator.
+$\frac{d}{dt}\boldsymbol{x}(t)=\frac{\mathrm{i}}{\hbar}[\boldsymbol{H},\boldsymbol{x}(t)],$ (2)
+where $\boldsymbol{x}(t)=\boldsymbol{U}_{t}\boldsymbol{x}(0)\boldsymbol{U}^{\dagger}_{t}$, $\boldsymbol{U}_{t}=\exp(\mathrm{i}\boldsymbol{H}t/\hbar)$ being the unitary time evolution operator.
 Since $\boldsymbol{U}_{t}$ commutes with $\boldsymbol{H}$, and $\boldsymbol{x}$ commutes with $V(\boldsymbol{x})$, the above equation yields
@@
 Similarly, the Heisenberg equation of motion for the momentum operator can be written as
-$\frac{d}{dt}\boldsymbol{p}(t)=\frac{\iota}{\hbar}[\boldsymbol{H},\boldsymbol{p}(t)]=\frac{\iota}{\hbar}\boldsymbol{U}_{t}[V(\boldsymbol{x}),\boldsymbol{p}(0)]\boldsymbol{U}^{\dagger}_{t}.$ (4)
+$\frac{d}{dt}\boldsymbol{p}(t)=\frac{\mathrm{i}}{\hbar}[\boldsymbol{H},\boldsymbol{p}(t)]=\frac{\mathrm{i}}{\hbar}\boldsymbol{U}_{t}[V(\boldsymbol{x}),\boldsymbol{p}(0)]\boldsymbol{U}^{\dagger}_{t}.$ (4)
 Now let us assume that the potential $V$ is an analytical function of $x$ and it can be expanded in a Taylor series, say, around $x_{0}$
@@
 $[V(\boldsymbol{x}),\boldsymbol{p}(0)]$ $=$ $\sum_{n=0}^{\infty}\frac{V^{(n)}(x_{0})}{n!}[(\boldsymbol{x}-x_{0})^{n},\boldsymbol{p}]$ (6)
-$=$ $\iota\hbar\sum_{n=0}^{\infty}\frac{V^{(n)}(x_{0})}{n!}n(\boldsymbol{x}-x_{0})^{n-1}$
+$=$ $\mathrm{i}\hbar\sum_{n=0}^{\infty}\frac{V^{(n)}(x_{0})}{n!}n(\boldsymbol{x}-x_{0})^{n-1}$
 $=$ $i\hbar V^{\prime}(\boldsymbol{x}),$
-where we have used the well known commutation relation $\boldsymbol{x}^{n},\boldsymbol{p}]=\iota\hbar n\boldsymbol{x}^{n-1}$.
+where we have used the well known commutation relation $\boldsymbol{x}^{n},\boldsymbol{p}]=\mathrm{i}\hbar n\boldsymbol{x}^{n-1}$.
 Then the Heisenberg equation of motion for $\boldsymbol{p}$ (4) assumes the form
@@
 which is the same as Newton’s equation for a Harmonic oscillator.
-It is easy to see that generalization to three dimensions and many particles is straightforward, simply because of observables of different particles commute with each others.
+It is easy to see that generalization to three dimensions and many particles is straightforward, simply because observables of different particles commute with each other.
 As an example, we focus on a small mass $m$ that is orbiting around a larger mass $M$.
@@
 The velocity operator is defined via the Heisenberg equation of motion:
-$\boldsymbol{v}_{i}=\frac{d\boldsymbol{r}_{i}}{dt}=\frac{\iota}{\hbar}\left[\boldsymbol{H},\boldsymbol{r}_{i}\right],$ (12)
+$\boldsymbol{v}_{i}=\frac{d\boldsymbol{r}_{i}}{dt}=\frac{\mathrm{i}}{\hbar}\left[\boldsymbol{H},\boldsymbol{r}_{i}\right],$ (12)
 where the indices $i=1,2,3$ represent the $x,y,z$ component of the relevant quantity, respectively. Using the canonical commutation relations
-$\boldsymbol{r}_{i},\boldsymbol{p}_{j}]=\iota\hbar\,\delta_{ij},\qquad[\boldsymbol{r}_{i},A_{j}(\boldsymbol{r})]=0,$ (13)
+$\boldsymbol{r}_{i},\boldsymbol{p}_{j}]=\mathrm{i}\hbar\,\delta_{ij},\qquad[\boldsymbol{r}_{i},A_{j}(\boldsymbol{r})]=0,$ (13)
 we compute
@@
 $=\frac{1}{2m}\left((\boldsymbol{p}_{j}-qA_{j})[\boldsymbol{p}_{j},\boldsymbol{r}_{i}]+[\boldsymbol{p}_{j},\boldsymbol{r}_{i}](\boldsymbol{p}_{j}-qA_{j})\right)$
-$=-\frac{\iota\hbar}{m}(\boldsymbol{p}_{i}-qA_{i}).$ (14)
+$=-\frac{\mathrm{i}\hbar}{m}(\boldsymbol{p}_{i}-qA_{i}).$ (14)
 Therefore, the velocity operator is
@@
 The Heisenberg equation for $\boldsymbol{\pi}_{i}$ reads
-$\frac{d\boldsymbol{\pi}_{i}}{dt}=\frac{\iota}{\hbar}\left[\boldsymbol{H},\boldsymbol{\pi}_{i}\right].$ (17)
+$\frac{d\boldsymbol{\pi}_{i}}{dt}=\frac{\mathrm{i}}{\hbar}\left[\boldsymbol{H},\boldsymbol{\pi}_{i}\right].$ (17)
 Since
@@
 we obtain
-$\boldsymbol{H},\boldsymbol{\pi}_{i}]=\frac{1}{2m}\left(\boldsymbol{\pi}_{j}[\boldsymbol{\pi}_{j},\boldsymbol{\pi}_{i}]+[\boldsymbol{\pi}_{j},\boldsymbol{\pi}_{i}]\boldsymbol{\pi}_{j}\right)-\iota\hbar q\mathcal{E}_{i}.$ (19)
+$\boldsymbol{H},\boldsymbol{\pi}_{i}]=\frac{1}{2m}\left(\boldsymbol{\pi}_{j}[\boldsymbol{\pi}_{j},\boldsymbol{\pi}_{i}]+[\boldsymbol{\pi}_{j},\boldsymbol{\pi}_{i}]\boldsymbol{\pi}_{j}\right)-\mathrm{i}\hbar q\mathcal{E}_{i}.$ (19)
 Using
-$\boldsymbol{p}_{i},f(\boldsymbol{r})]=-\iota\hbar\,\partial_{i}f(\boldsymbol{r}),$ (20)
+$\boldsymbol{p}_{i},f(\boldsymbol{r})]=-\mathrm{i}\hbar\,\partial_{i}f(\boldsymbol{r}),$ (20)
 we compute
@@
 $=-q[\boldsymbol{p}_{i},A_{j}]-q[A_{i},\boldsymbol{p}_{j}]$
-$=\iota\hbar q(\partial_{i}A_{j}-\partial_{j}A_{i})$
-$=\iota\hbar q\sum_{k=1}^{3}\epsilon_{ijk}B_{k},$ (21)
+$=\mathrm{i}\hbar q\left(\frac{\partial A_{j}}{\partial r_{i}}-\frac{\partial A_{i}}{\partial r_{j}}\right)$
+$=\mathrm{i}\hbar q\sum_{k=1}^{3}\epsilon_{ijk}B_{k},$ (21)
 where $\epsilon_{ijk}$ is the Levi-Civita epsilon. Substituting into the Hamiltonian commutator,
-$\boldsymbol{H},\boldsymbol{\pi}_{i}]=\frac{\iota\hbar q}{m}\sum_{j,k}\epsilon_{ijk}\boldsymbol{\pi}_{j}B_{k}-\iota\hbar q\mathcal{E}_{i}.$ (22)
+$\boldsymbol{H},\boldsymbol{\pi}_{i}]=\frac{\mathrm{i}\hbar q}{m}\sum_{j,k}\epsilon_{ijk}\boldsymbol{\pi}_{j}B_{k}-\mathrm{i}\hbar q\mathcal{E}_{i}.$ (22)
 Therefore,
-$m\frac{d\boldsymbol{v}_{i}}{dt}$ $=\frac{\iota}{\hbar}[\boldsymbol{H},\boldsymbol{\pi}_{i}]$ (23)
+$m\frac{d\boldsymbol{v}_{i}}{dt}$ $=\frac{\mathrm{i}}{\hbar}[\boldsymbol{H},\boldsymbol{\pi}_{i}]$ (23)
 $=q\sum_{j,k}\epsilon_{ijk}\boldsymbol{v}_{j}B_{k}+q\mathcal{E}_{i},$ (24)
@@
 Remarkably, the Lorentz force law emerges exactly from operator algebra in quantum mechanics.
-IV Ehrenfest theorem
+IV Rotating frame of reference
+Now that we have seen the Lorentz force law emerging in quantum mechanics, we wish to look for other situations where velocity dependent forces arise.
+It is well known in classical mechanics that if we use a rotating frame of reference, Coriolis force and centrifugal force arise as pseudo forces.
+Let us explore this situation in quantum mechanics.
+Consider a reference frame $\text{S}^{\prime}$ that rotates with a constant angular velocity $\omega$ around the origin of coordinates in the $xy$ plane of frame S.
+The Hamiltonian as seen by an observer in the rotating frame is given by pseudoforces
+$\boldsymbol{H}=\frac{1}{2m}\Big[(\boldsymbol{p}_{1}+m\omega\boldsymbol{r}_{2})^{2}+(\boldsymbol{p}_{2}-m\omega\boldsymbol{r}_{1})^{2}\Big]-m\omega^{2}(\boldsymbol{r}_{1}^{2}+\boldsymbol{r}_{2}^{2}),$ (26)
+where we follow the notions of the previous section. The Heisenberg equations for positions are given by
+$\frac{d\boldsymbol{r}_{1}}{dt}=\frac{1}{m}(\boldsymbol{p}_{1}+m\omega\boldsymbol{r}_{2}),~~~~~~\frac{d\boldsymbol{r}_{2}}{dt}=\frac{1}{m}(\boldsymbol{p}_{2}-m\omega\boldsymbol{r}_{1}).$
+Denoting $\boldsymbol{v}_{i}={d\boldsymbol{r}_{i}}/{dt}$ given by the above equations, we can write the Heisenberg equations for these velocities
+$\frac{d\boldsymbol{v}_{i}}{dt}=\frac{\mathrm{i}}{\hbar}[H,\boldsymbol{v}_{i}].$ (28)
+Evaluating the various commutators involved, we obtain
+$\frac{d\boldsymbol{v}_{1}}{dt}=2\omega\boldsymbol{v}_{2}+\omega^{2}\boldsymbol{r}_{1},~~~~~~\frac{d\boldsymbol{v}_{2}}{dt}=-2\omega\boldsymbol{v}_{1}+\omega^{2}\boldsymbol{r}_{2}.$
+Using $\vec{\omega}=\omega\,\hat{z}$, these equations combine to
+$m\frac{d^{2}\vec{\boldsymbol{r}}}{dt^{2}}=-2m\,\vec{\omega}\times\vec{\boldsymbol{v}}-m\vec{\omega}\times(\vec{\omega}\times\vec{\boldsymbol{r}}).$ (30)
+The equation presented above is fully quantum mechanical, and is derived from the Heisenberg equations of motion.
+On the right hand side one can identify $m\vec{\omega}\times(\vec{\omega}\times\vec{\boldsymbol{r}})$ with the centrifugal force and $2m\,\vec{\omega}\times\vec{\boldsymbol{v}}$ with the Coriolis force.
+It is identical to Newton’s second law of motion in a rotating frame of reference.
+V Ehrenfest theorem
 Till now we have studied the dynamics of quantum observables at the operator level.
@@
 where the angular brackets denote the expectation value.
-This is the familiar equation denoting Ehrefest theorem Ehrenfest .
+This is the familiar equation denoting Ehrenfest theorem Ehrenfest .
 It shows that the expection values of position and momentum approximately follow Newton’s equations of motion.
@@
 For a harmonic oscillator, since $V^{\prime}(\boldsymbol{x})$ is linear in $\boldsymbol{x}$, $\langle V^{\prime}({\boldsymbol{x}})\rangle=V^{\prime}(\langle{\boldsymbol{x}}\rangle)$, and the position and momentum expectation values exactly follow classical dynamics, for all quantum states.
-V Conclusions
+It is straightforward to see that since eqns.
+(25) and (30) are linear in the observables, both of them would lead to the expectation values of observables exactly following classical dynamics.
+VI Generalized treatment
+Now that we are clear that quantum and classical dynamics is governed by the same equations in a wide variety of situations, we wish to demonstrate it in a general scenario.
+Consider a system described by the Hamiltonian as a function of the generalized coordinates and generalized momenta, $\boldsymbol{H}(\boldsymbol{q}_{i},\boldsymbol{p}_{i})$.
+Heisenberg equations for each of the generalized coordinates and momenta are given by
+$\frac{d\boldsymbol{q}_{j}}{dt}=\frac{\mathrm{i}}{\hbar}[\boldsymbol{H}(\boldsymbol{q}_{i},\boldsymbol{p}_{i}),\boldsymbol{q}_{j}],~~~~~~\frac{d\boldsymbol{p}_{j}}{dt}=\frac{\mathrm{i}}{\hbar}[\boldsymbol{H}(\boldsymbol{q}_{i},\boldsymbol{p}_{i}),\boldsymbol{p}_{j}].$ (33)
+Assuming that $\boldsymbol{H}$ can be expanded in a Taylor series in each generalized coordinate and momentum, we can write
+$\boldsymbol{H}$ $=$ $\sum_{n=0}^{\infty}\frac{(\boldsymbol{p}_{j}-a_{j})^{n}}{n!}\left.\frac{\partial^{n}\boldsymbol{H}}{\partial p_{j}^{n}}\right|_{p_{j}=a_{j}}$
+$\boldsymbol{H}$ $=$ $\sum_{n=0}^{\infty}\frac{(\boldsymbol{q}_{j}-b_{j})^{n}}{n!}\left.\frac{\partial^{n}\boldsymbol{H}}{\partial q_{j}^{n}}\right|_{q_{j}=b_{j}},$ (34)
+where a derivative like $\frac{\partial\boldsymbol{H}}{\partial p_{i}}$ should be understood as taking the derivative of $\boldsymbol{H}$ with respect to a variable $q_{j}$ by treating it as a classical variable, and then replacing it by the quantum observable in the result.
+Now we obtain
+$\big[\boldsymbol{H}(\boldsymbol{q}_{i},\boldsymbol{p}_{i}),\boldsymbol{q}_{j}\big]$ $=$ $-\,\mathrm{i}\hbar\sum_{n=0}^{\infty}\frac{n(\boldsymbol{p}_{j}-a_{j})^{n-1}}{n!}\left.\frac{\partial^{n}\boldsymbol{H}}{\partial p_{j}^{n}}\right|_{p_{j}=a_{j}},$
+$\big[\boldsymbol{H}(\boldsymbol{q}_{i},\boldsymbol{p}_{i}),\boldsymbol{p}_{j}\big]$ $=$ $\,\mathrm{i}\hbar\sum_{n=0}^{\infty}\frac{n(\boldsymbol{q}_{j}-b_{j})^{n-1}}{n!}\left.\frac{\partial^{n}\boldsymbol{H}}{\partial q_{j}^{n}}\right|_{q_{j}=b_{j}}.$ (35)
+Here, we use the identities
+$\big[(\boldsymbol{p}_{i}-a_{i})^{n},\boldsymbol{q}_{j}\big]$ $=$ $-\,\mathrm{i}\hbar\,n\,(\boldsymbol{p}_{i}-a_{i})^{n-1}\,\delta_{ij},$
+$\big[(\boldsymbol{q}_{i}-b_{i})^{n},\boldsymbol{p}_{j}\big]$ $=$ $\;\,\mathrm{i}\hbar\,n\,(\boldsymbol{q}_{i}-b_{i})^{n-1}\,\delta_{ij}.$ (36)
+Substituting (35) in (33), we obtain
+$\frac{d\boldsymbol{q}_{j}}{dt}$ $=$ $\frac{\partial}{\partial\boldsymbol{p}_{j}}\sum_{n=0}^{\infty}\frac{(\boldsymbol{p}_{j}-a_{j})^{n}}{n!}\left.\frac{\partial^{n}\boldsymbol{H}}{\partial p_{j}^{n}}\right|_{p_{j}=a_{j}},$
+$\frac{d\boldsymbol{p}_{j}}{dt}$ $=$ $-\frac{\partial}{\partial\boldsymbol{q}_{j}}\sum_{n=0}^{\infty}\frac{(\boldsymbol{q}_{j}-b_{j})^{n}}{n!}\left.\frac{\partial^{n}\boldsymbol{H}}{\partial q_{j}^{n}}\right|_{q_{j}=b_{j}}.$ (37)
+It should be kept in mind that $\left.\frac{\partial^{n}\boldsymbol{H}}{\partial p_{j}^{n}}\right|_{p_{j}=a_{j}}$ is still a function of all the $\boldsymbol{q}_{i}$s and all the $\boldsymbol{p}_{i}$s except $\boldsymbol{p}_{j}$.
+Similarly, $\left.\frac{\partial^{n}\boldsymbol{H}}{\partial q_{j}^{n}}\right|_{q_{j}=b_{j}}$ is a function of all the $\boldsymbol{p}_{i}$s and all the $\boldsymbol{q}_{i}$s except $\boldsymbol{q}_{j}$.
+The above can then be written in a more compact form:
+$\frac{d\boldsymbol{q}_{j}}{dt}$ $=$ $\frac{\partial\boldsymbol{H}}{\partial\boldsymbol{p}_{j}},~~~~~~\frac{d\boldsymbol{p}_{j}}{dt}=-\frac{\partial\boldsymbol{H}}{\partial\boldsymbol{q}_{j}}.$ (38)
+Remarkably eqns.
+(38) are a modified form of the Heisenberg equations, but have the exact form of the classical Hamilton’s equations of motion, if one were to substitute the operators by the corresponding classical variables.
+VII Conclusions
 In this work, we have demonstrated that the dynamical equations governing classical and quantum mechanics possess a deeper and more exact correspondence than is usually emphasized.
@@
 For a wide class of systems, ranging from particles moving in arbitrary analytic potentials to charged particles in external electromagnetic fields, the Heisenberg equations reproduce the familiar classical laws of motion, including Newton’s second law and the Lorentz force equation.
+In the general treatment, the Heisenberg equations reproduce the Hamilton’s equations of motion.
 The absence of $\hbar$ in these operator equations highlights that the distinction between classical and quantum dynamics does not arise from the form of the equations themselves, but rather from the non-commuting nature of quantum observables and the role of the quantum state.
@@
 82-27/2019 (SA-III).
-TQ acknowledges useful discussions with Bibhash Paul.
+TQ acknowledges useful discussions with Bibhash Paul, T.P.
+Singh and Pankaj Sharan.
 Author declarations